# The Strong Cosmic Censorship Conjecture

### v1

@@ 5.2 From Weak Cosmic Censorship to Strong Cosmic Censorship
While Theorem 5.2 only provides $C^{2}$ inextendibility, the spacelike nature of the terminal MGHD boundary $\mathcal{S}=\{r=0\}$ (and Theorem 3.2) leads to conjecturing that a generic spherically-symmetric solution is $C^{0}$ inextendible. This expectation is reinforced by the fact that $\mathcal{S}=\{r=0\}$ is asymptotically Schwarzschild towards $i^{+}$ as proven in [102] (see also [103, 104]). For the more sophisticated Einstein–Maxwell-charged-scalar-field model (4.4) which additionally allows to study Cauchy horizons emanating from $i^{+}$ (as we explained in Section 4.3), the global picture of spherically-symmetric gravitational collapse is more subtle and several aspects will be discussed in Section 5.3 and Section 9. We emphasize, however, that the Christodoulou argument showing that $\mathcal{CH}_{\Gamma}=\emptyset$ generically is local near $\Gamma$, and thus, independent of the dynamics near $i^{+}$. We may thus speculate [13, 87] that $\mathcal{CH}_{\Gamma}=\emptyset$ for generic spherically-symmetric solutions of the Einstein–Maxwell-charged-scalar-field model (4.4) (see [97, 105] for recent results in this direction).
@@ 5.3 The breakdown of weak null singularities
It is important to point out that the examples of [106] with the Penrose diagram of Figure 8 have a non-singular Cauchy horizon $\mathcal{CH}^{+}$ (consistently with Theorem 5.3), a situation which is conjectured to be non-generic. Indeed, in Theorem 4.7, we recall that $\mathcal{CH}^{+}$ was shown to be weakly singular under decay assumptions on the event horizon that are conjectured to be generic. The main mechanism in Theorem 5.3 – the “breakdown of the weak null singularity” $\mathcal{CH}^{+}$ – is proved by a contradiction argument, using the focusing properties of the Einstein equations (notably the null constraint, so-called Raychaudhuri, equations) and taking advantage of the strength of the singularity at $\mathcal{CH}^{+}$ combined with the spacetime geometry near the central axis $\Gamma$. In view of the breakdown of $\mathcal{CH}^{+}$, there must be another (presumably spacelike, see Figure 9) component of the terminal boundary; we will discuss this aspect in Section 9 below.
@@ 5.4 The spacelike portion of the terminal boundary
Assume the spacetime in Theorem 5.3 does not possess a locally naked singularity, i.e., $\mathcal{CH}_{\Gamma}=\emptyset$, and denote $\mathcal{B}$ its terminal boundary. Then $\mathcal{B}$ contains a spacelike singularity $\mathcal{S}\neq\emptyset$ intersecting $\mathcal{CH}^{+}\subset\{v=\infty\}$ and the metric near $\mathcal{CH}^{+}\cap\mathcal{S}$ is approximated by a Kasner metric of $v$-dependent Kasner exponents $(1-2p(u,v),p(u,v),p(u,v))$, where $p(u,v)\approx\frac{C}{v}$ as $v\rightarrow\infty$ for some $C>0$.
Moreover, in the companion paper [108], concrete examples of one-ended spacetimes satisfying the assumptions of Theorem 5.4 are constructed.
@@
Theorem 5.4 provides the first examples of a null-spacelike singularity transition [^21] Theorem 5.4 is, in fact, the corollary of a more general local result in [95]. In particular, the Kasner-like quantitative estimates of Theorem 5.4 also apply to the two-ended black holes with $\mathcal{S}\neq\emptyset$ as in the rightmost diagram of Figure 7, see [108]. in a black hole, together with the first construction a Kasner-like singularity with degenerating exponents (see [107] for a very general construction of Kasner-type spacetimes, which however assumes the Kasner exponents to be non-degenerating and thus does not cover the situation of Theorem 5.4). Lastly, one may speculate that the results of Theorem 5.4 can be suitably generalized [^22] The landmark result of Fournodavlos–Rodninanski–Speck [92] indeed proves that Kasner metrics with constant positive exponents are stable to small perturbations. The Kasner exponents in Theorem 5.4 are also positive, thus the spacetime is expected to be stable against non-spherical perturbations. However, the Kasner exponents in Theorem 5.4 are non-constant and degenerate as $v\rightarrow+\infty$, and therefore the techniques of [92] do not apply. to non spherically-symmetric spacetimes solving the Einstein-scalar-field system (2.4), (3.7). In summary, it is conjectured that the geometry of generic (non spherically-symmetric) Einstein-scalar-field solutions is as in Figure 9. The situation for the Einstein vacuum equations, however, is more uncertain in that Kasner-type singularities are not expected to be stable, see [107, 109].
@@ 6.1 Linear wave equations on a stationary black hole and Price’s law
While (6.1) is often considered as a linear toy model for the Einstein equations (2.4) in a vacuum, the equations of linearized gravity around the Kerr metric (3.5) give rise to the celebrated Teukolsky equations [122] which are spin-weighted analogues of (6.1) and have the following form on (sub-extremal) Kerr spacetime (3.5), where the spin $s$ takes the value {±$\{\pm 2\}$.
@@ 6.4 Wave equations on dynamical black holes and black hole stability
Given any asymptotically flat, spherically symmetric, admissible data on ×$\mathbb{R}\times\mathbb{S}^{2}$, the MGHD $(\mathcal{M},g,\phi)$ for (4.2) converges to a Reissner–Nordström black hole at the following rate: there exists $D\in\mathbb{R}$, $\epsilon>0$, $C\in\mathbb{R}$ such that
@@ 7.1 Exceptional dynamical black holes arising from finite co-dimension initial data
Asymptotically Schwarzschildean black holes. Within spherical symmetry, black hole solutions of (3.7) with localized initial data are all asymptotically Schwarzschildean (this follows from [74]) and the MGHD features a spacelike singularity $\mathcal{S}=\{r=0\}$, as discussed in Section 5.2. For spherically-symmetric models with charge, in which asymptotically Schwarzschildean black holes are non-generic, it is shown in the author’s [159] that asymptotically Schwarzschildean spherically-symmetric solutions of (4.4) have no Cauchy horizon ($\mathcal{CH}^{+}=\emptyset$) and the MGHD terminal boundary is foliated by spheres of area-radius $r=0$. For the Einstein vacuum equations, the result of Dafermos–Holzegel–Rodnianski–Taylor [151] shows that asymptotically Schwarzschildean black holes arise from a co-dimension three subset of initial data close to Schwarzschild. For these asymptotically Schwarzschildean black holes, the spacetime supremum of Kretschmann curvature blows up [151, 22] in the black hole interior, which suggests that the MGHD is future-inextendible. However, the causal nature (null or spacelike) of the singularity, and quantitative estimates in its vicinity, however, remain open, even within spherical symmetry.
@@ 7.2 Spatially-compact spacetimes, cosmological analogues of Cosmic Censorship
Strong Cosmic Censorship in the vicinity of Kasner spacetimes. Recall the Kasner metric (4.10) from Section 4.7 is an anisotropic spatially-homogeneous spacetime with compact spatial topology $\mathbb{T}^{3}$ and a spacelike singularity at $\{t=0\}$. Its perturbations for the Einstein vacuum equations may have very intricate dynamics (BKL heuristics [93, 94, 173]).However, in the presence of a scalar field (i.e., for solutions of (2.4), (3.7)) or within $\mathbb{U}(1)$-polarized solutions of the Einstein vacuum equations, there exist Kasner metrics (4.10) with positive exponents $(p_{1},p_{2},p_{3})$. The work of Fournodavlos–Rodnianski–Speck [92] proves the stability of such solutions, and one can thus infer a version of Strong Cosmic Censorship in a vicinity of these Kasner spacetimes for the Einstein-scalar-field equations (or the vacuum Einstein equations in $\mathbb{U}(1)$-polarized symmetry). The results of [92] were recently generalized by Groeninger–Peterson–Ringström [174] proving the stability of a larger class of so-called “quiescent singularities” that includes Kasner metrics (4.10) with positive exponents $p_{i}$. Thus, an analogous statement of Strong Cosmic Censorship can be obtained in the vicinity of the larger class of initial data considered in [174].
@@ 8 Summary and future directions
Strong Cosmic Censorship research stands as a vibrant frontier, pushing the boundaries of both general relativity and mathematical analysis. Significant progress has been made, yet captivating challenges remain. As we delve deeper into this fascinating domain, we stand to unlock a wealth of knowledge about the nature of gravity and the enigmatic realm of black holes.

### v2

@@ 5.2 From Weak Cosmic Censorship to Strong Cosmic Censorship
While Theorem 5.2 only provides $C^{2}$ inextendibility, the spacelike nature of the terminal MGHD boundary $\mathcal{S}=\{r=0\}$ (and Theorem 3.2) leads to conjecturing that a generic spherically-symmetric solution is $C^{0}$ inextendible. This expectation is reinforced by the fact that $\mathcal{S}=\{r=0\}$ is asymptotically Schwarzschild towards $i^{+}$ as proven in [104] (see also [105, 106]). For the more sophisticated Einstein–Maxwell-charged-scalar-field model (4.4) which additionally allows to study Cauchy horizons emanating from $i^{+}$ (as we explained in Section 4.3), the global picture of spherically-symmetric gravitational collapse is more subtle and several aspects will be discussed in Section 5.3 and Section 5.4. We emphasize, however, that the Christodoulou argument showing that $\mathcal{CH}_{\Gamma}=\emptyset$ generically is local near $\Gamma$, and thus, independent of the dynamics near $i^{+}$. We may thus speculate [13, 87] that $\mathcal{CH}_{\Gamma}=\emptyset$ for generic spherically-symmetric solutions of the Einstein–Maxwell-charged-scalar-field model (4.4).
@@ 5.3 The breakdown of weak null singularities
It is important to point out that the examples of [107] with the Penrose diagram of Figure 8 have a non-singular Cauchy horizon $\mathcal{CH}^{+}$ (consistently with Theorem 5.3), a situation which is conjectured to be non-generic. Indeed, in Theorem 4.7, we recall that $\mathcal{CH}^{+}$ was shown to be weakly singular under decay assumptions on the event horizon that are conjectured to be generic. The main mechanism in Theorem 5.3 – the “breakdown of the weak null singularity” $\mathcal{CH}^{+}$ – is proved by a contradiction argument, using the focusing properties of the Einstein equations (notably the null constraint, so-called Raychaudhuri, equations) and taking advantage of the strength of the singularity at $\mathcal{CH}^{+}$ combined with the spacetime geometry near the central axis $\Gamma$. In view of the breakdown of $\mathcal{CH}^{+}$, there must be another (presumably spacelike, see Figure 9) component of the terminal boundary; we will discuss this aspect in Section 5.4 below.
@@ 5.4 The spacelike portion of the terminal boundary
Assume the spacetime in Theorem 5.3 does not possess a locally naked singularity, i.e., $\mathcal{CH}_{\Gamma}=\emptyset$, and denote $\mathcal{B}$ its terminal boundary. Then $\mathcal{B}$ contains a spacelike singularity $\mathcal{S}\neq\emptyset$ intersecting $\mathcal{CH}^{+}\subset\{v=\infty\}$ and the metric near $\mathcal{CH}^{+}\cap\mathcal{S}$ is approximated by a Kasner metric of $v$-dependent Kasner exponents (12p(u,v),p(u,v),p(u,v))(1-2p(u,v),p(u,v),p(u,v)), where $p(u,v)\approx\frac{C}{v}$ as $v\rightarrow\infty$ for some $C>0$.
While Theorem 5.4 in a local result in the black hole interior, the companion paper [97] provides concrete global examples of one-ended asymptotically flat spacetimes satisfying the assumptions of Theorem 5.4:
Theorem 5.5 ([97]).
There exist examples of spherically symmetric one-ended asymptotically flat black hole solution of (4.4) such that the terminal boundary has two non-empty components: a weakly singular null Cauchy horizon $\mathcal{CH}^{+}\neq$$\emptyset$, and a crushing singularity $\mathcal{S}=\{r=0\}$. Moreover, $\mathcal{S}$ is spacelike near $\mathcal{CH}^{+}\cap\mathcal{S}$ and obeys the Kasner estimates of Theorem 5.4.
@@
Theorem 5.4 provides the first examples of a null-spacelike singularity transition [^20] Theorem 5.4 is, in fact, the corollary of a more general local result in [98]. In particular, the Kasner-like quantitative estimates of Theorem 5.4 also apply to the two-ended black holes with $\mathcal{S}\neq\emptyset$ as in the rightmost diagram of Figure 7, see [97]. in a black hole, together with the first construction a Kasner-like singularity with degenerating exponents (see [108] for a very general construction of Kasner-type spacetimes, which however assumes the Kasner exponents to be non-degenerating and thus does not cover the situation of Theorem 5.4). Lastly, one may speculate that the results of Theorem 5.4 can be suitably generalized [^21] The landmark result of Fournodavlos–Rodnianski–Speck [92] indeed proves that Kasner metrics with constant positive exponents are stable to small perturbations. The Kasner exponents in Theorem 5.4 are also positive, thus the spacetime is expected to be stable against non-spherical perturbations. However, the Kasner exponents in Theorem 5.4 are non-constant and degenerate as $v\rightarrow+\infty$, and therefore the techniques of [92] do not apply. to non spherically-symmetric spacetimes solving the Einstein-scalar-field system (2.4), (3.7). In summary, it is conjectured that the geometry of generic (non spherically-symmetric) Einstein-scalar-field solutions is as in Figure 9. The situation for the Einstein vacuum equations, however, is more uncertain in that Kasner-type singularities are not expected to be stable, see [108, 109].
@@ 6.1 Linear wave equations on a stationary black hole and Price’s law
While (6.1) is often considered as a linear toy model for the Einstein equations (2.4) in a vacuum, the equations of linearized gravity around the Kerr metric (3.5) give rise to the celebrated Teukolsky equations [122] which are spin-weighted analogues of (6.1) and have the following form on (sub-extremal) Kerr spacetime (3.5), where the spin $s$ takes the value $\{\pm 2\}$.
@@ 6.4 Wave equations on dynamical black holes and black hole stability
Given any asymptotically flat, spherically symmetric, admissible data on $\mathbb{R}\times\mathbb{S}^{2}$, the MGHD $(\mathcal{M},g,\phi)$ for (4.2) converges to a Reissner–Nordström black hole at the following rate: there exists $D\in\mathbb{R}$, $\epsilon>0$, $C\in\mathbb{R}$ such that
@@ 7.1 Exceptional dynamical black holes arising from finite co-dimension initial data
Asymptotically Schwarzschildean black holes. Within spherical symmetry, black hole solutions of (3.7) with localized initial data are all asymptotically Schwarzschildean (this follows from [74]) and the MGHD features a spacelike singularity $\mathcal{S}=\{r=0\}$, as discussed in Section 5.2. For spherically-symmetric models with charge, in which asymptotically Schwarzschildean black holes are non-generic, it is shown in the author’s [159] that asymptotically Schwarzschildean spherically-symmetric solutions of (4.4) have no Cauchy horizon ($\mathcal{CH}^{+}=\emptyset$) and the MGHD terminal boundary is foliated by spheres of area-radius $r=0$. For the Einstein vacuum equations, the result of Dafermos–Holzegel–Rodnianski–Taylor [151] shows that asymptotically Schwarzschildean black holes arise from a co-dimension three subset of initial data close to Schwarzschild. For these asymptotically Schwarzschildean black holes, the spacetime supremum of Kretschmann curvature blows up [151, 22] in the black hole interior, which suggests that the MGHD is future-inextendible. However, the causal nature (null or spacelike) of the singularity, and quantitative estimates in its vicinity remain open, even within spherical symmetry.
@@ 7.2 Spatially-compact spacetimes, cosmological analogues of Cosmic Censorship
Strong Cosmic Censorship in the vicinity of Kasner spacetimes. Recall the Kasner metric (4.11) from Section 4.7 is an anisotropic spatially-homogeneous spacetime with compact spatial topology $\mathbb{T}^{3}$ and a spacelike singularity at $\{t=0\}$. Its perturbations for the Einstein vacuum equations may have very intricate dynamics (BKL heuristics [93, 94, 173]). However, the dynamics is expected to admit stable fixed points represented by Kasner metrics of the form (4.11) either the presence of a scalar field (i.e., for solutions of (2.4), (3.7)) or within $\mathbb{U}(1)$-polarized solutions of the Einstein vacuum equations. The work of Fournodavlos–Rodnianski–Speck [92] proves the stability of such solutions, and one can thus infer a version of Strong Cosmic Censorship in a vicinity of these Kasner spacetimes for the Einstein-scalar-field equations (or the vacuum Einstein equations in $\mathbb{U}(1)$-polarized symmetry). The results of [92] were recently generalized by Groeninger–Peterson–Ringström [174] proving the stability of a larger class of so-called “quiescent singularities” that contains the class considered in [92]. Thus, an analogous statement of Strong Cosmic Censorship can be obtained in the vicinity of the larger class of initial data considered in [174].
@@ 8 Summary and future directions
Strong Cosmic Censorship research stands as a vibrant frontier, pushing the boundaries of both General Relativity and mathematical analysis. Significant progress has been made, yet captivating challenges remain. As we delve deeper into this captivating domain, we stand to unlock a wealth of knowledge about the nature of gravity and the enigmatic realm of black holes.